OpenStudy (bethgrace96):

IF A CITYS POP GROWTH RATE IS 7% PER YEAR (COMPOUNDED ANUALLY) , HOW LONG WILL IT TAKE THE CITYS POP TO DOUBLE? A) 3.86 yrs B)9.90 yrs C)10.24 yrs D) 26 yrs

OpenStudy (kropot72):

Let P be the starting population. Then we can use the following equation: \[2P=P(1+0.07)^{t}\] which simplifies to \[2=1.07^{t}\] Taking natural logs of both sides we get: \[\ln 2=t\ \ln 1.07\] from which can can find the value of t: \[t=\frac{\ln 2}{\ln 1.07}=you\ can\ calculate\]

OpenStudy (kropot72):

@bethgrace96 Are you there?

OpenStudy (bethgrace96):

thanks so much!!

OpenStudy (kropot72):

You're welcome :)